OpenStudy (anonymous):

Find dy/dx by implicit differentiation. 7y sin(x^2) = 9x sin(y^2)

OpenStudy (anonymous):

take the derivative of every term, from left to right. use the chain rule (including y') and the product rule well...

OpenStudy (anonymous):

Left side...

OpenStudy (anonymous):

would dy/dx of 7y be 7yprime or yprime

OpenStudy (anonymous):

7 y'

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

ready for left side?

OpenStudy (anonymous):

do you mean the right?

OpenStudy (anonymous):

No, I meant left, to check in. Did you get: 7 y' sin(x^2) + 7 y cos(x^2) 2x

OpenStudy (anonymous):

yes :D

OpenStudy (anonymous):

Great, let me know if you have any questions about the right or solving for y'

OpenStudy (anonymous):

would it be like (9x)(cos(2y)(yprime)+(9)(sin(y^2)

OpenStudy (anonymous):

(9x)(cos(2y)(yprime)+(9)(sin(y^2) Chain rule error. When you do the f'(g(x)) part of the chain rule, re-write the inner function - the y^2 without taking its derivative yet (next post I'll show you)

OpenStudy (anonymous):

9x cos(y^2) 2y y' + 9 sin(y^2)

OpenStudy (anonymous):

okay i see where i messed up

OpenStudy (anonymous):

then i isolate \[y^l\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay im not sure if i did it right

OpenStudy (anonymous):

\[-9x[\sin(y^2)]+7y[\cos(x^2)] / 9x[\cos(y^2)\times2y]+7[\sin(x^2)]\]

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

wait its wrong

OpenStudy (anonymous):

I think I found two errors in the numerator and added parentheses: (−9 [sin(y2)]+7y 2x [cos(x2)])/(9x[cos(y2)×2y]+7[sin(x2)])

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

what you said